Spot FX rate is 3.600 PLN/USD. You can enter a 3-year forward contract with an exercise price of 4.0000 PLN/USD. The rates are fixed in the whole investment period and equal 10% pa. for Polish market and 5% for USA market, what is the theoretical price of this contract? Is arbitrage possible-if yes, explain why? If arbitrage is possible, what strategy should investor choose in order to get a riskless profit?
Answer #1

The theoretical price of forward contract = Spot Price * (1+ Polish Rate)3/(1+USA Market Rate)3 = 3.6*(1+10%)3/(1+5%)3 = 4.1392

Yes arbitrage is possible. The investor should buy a forward contract and after 3 years exercise the forward and sell it at higher exchange rate of 4.1392 So Profit in terms of PLN is 4.1392 - 4.000 = 0.1932 PLN

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well